# UAR-NVC: A Unified AutoRegressive Framework for Memory-Efficient Neural Video Compression

### v1

@@ IV-A2 Implementation Details
To achieve different bitrate points, we follow previous work to adjust the model size [16, 24, 25]. For small $p$, the model sizes are (0.3, 0.45, 0.6, 0.9) MB. For medium $p$, the model sizes are (0.6, 0.9, 1.8, 2.7) MB. For large $p$, the model sizes are (3.0, 4.5, 6.0, 9.0) MB. In addition, we set the rate-distortion trade-off parameter $\lambda$ to 5.0 for HiNeRV and 0.5 for HNeRV, as they have very different network architectures. We obtained these values using a fitting method similar to $\varepsilon$ described in Sec. III-E. The small, medium and large $p$ settings correspond to different learning rates and training epochs for HiNeRV: 5e-3/5e-3/2e-3 and 3k/1.5k/1.5k for I-model, and 2k/2k/1k for P-model. For HNeRV, the learning rates are adjusted to 2e-3/2e-3/1e-3. The papramter selection follows the setting in their original paper [25, 24]. Besides, we train for more epochs when $p$ is smaller, as fewer data points per epoch require longer training.
@@
    ffmpeg -pix_fmt yuv420p -s:v W×$\times$×H -i input.yuv -vframes N_e -c:v libx264 -preset veryslow -qp QP -g GOP output.mkv
@@
    ffmpeg -pix_fmt yuv420p -s:v W×$\times$×H -i Video.yuv -vframes Ne -c:v libx265 -preset veryslow -x265-params “qp=QP :keyint=GOP” output.mkv
@@ IV-A3 Evaluation Metrics
TABLE I: BD-rate (%) for PSNR/MS-SSIM. The anchor for each methods is x265 (veryslow).
Dataset UVG VVC ClassB
Metirc PSNR MS-SSIM PSNR MS-SSIM
GOP GOP6 GOP30 GOP120 GOP6 GOP30 GOP120 GOP5 GOP25 GOP100 GOP5 GOP25 GOP100
HNeRV (base) 358.00 225.81 157.71 254.53 140.22 108.40 397.38 413.73 343.36 332.77 389.09 279.90
HNeRV (ours) 189.30 113.17 113.04 234.61 135.19 76.12 291.54 250.67 193.87 251.86 238.46 169.31
HiNeRV (base) 217.16 102.15 65.89 37.28 4.56 -1.50 285.36 193.53 100.12 112.90 68.05 21.58
HiNeRV (ours) 20.43 -17.60 -34.29 -30.86 -51.29 -53.79 85.55 36.34 -18.79 3.01 -4.82 -50.57
@@
IV-B Rate-Distortion Performance
To evaluate the performance of our framework under different settings, we first conduct experiments on the UVG dataset. As shown in Fig. 7, both HiNeRV and HNeRV exhibit significant improvements when integrated with our proposed modules. Specifically, when using HNeRV (base) and HiNeRV (base) as anchors, and comparing the RD curve positions of HNeRV (ours) and HiNeRV (ours), we observe that the primary gain in HNeRV stems from bitrate savings, while the primary gain in HiNeRV arises from improved image quality. This improvement is attributed to HiNeRV’s more complex grid design, which enables it to capture finer details and fully leverage the potential of a more advanced entropy model with flexible bit allocation. Conversely, HNeRV’s architecture is optimized for efficiency, featuring a compact structure that reduces memory consumption and accelerates inference speed. This streamlined design also imposes a limit on its maximum reconstruction quality, even as it achieves substantial bitrate reductions with our two modules. When evaluated using the MS-SSIM metric, both models exhibit a trend similar to the analysis in PSNR. However, compared to x265 (veryslow) as the anchor, INR-based methods achieve significantly better in MS-SSIM relative to PSNR. This discrepancy arises because INR models encode video data in a more holistic manner, leading to superior structural fidelity. In contrast, traditional block-based encoders struggle to effectively capture the structural relationships between blocks.
Examining the performance gap across different $p$ settings, we observe that in smaller $p$ settings, the RD performance of INR-based model is inferior to that in larger $p$ settings. This is primarily because models with larger $p$ settings are more effective at reducing temporal redundancy. Consequently, choosing an appropriate $p$ value involves a trade-off dictated by resource constraints. In addition, we note that when $p=6$, HiNeRV (ours) performs less effectively than x265 (veryslow) at low bitrates but surpasses it at higher bitrates. This occurs because, when fitted with a smaller model, the HiNeRV structure has limited capacity to capture fine details, reducing its fitting ability. However, higher bitrate models alleviate this limitation, leading to improved overall performance. As $p$ increases, the performance of our method improves steadily and eventually surpasses that of x265 (veryslow), demonstrating the robust modeling capabilities of implicit representation models. Notably, $p=30$ represents a highly practical configuration, with encoding delays that remain acceptable in many real-world scenarios. In this setup, our RD performance exceeds that of x265 (veryslow), further validating the effectiveness of our framework. When $p=120$, HiNeRV (ours) significantly outperforms x265 (veryslow). This result demonstrates that our framework is not only suitable for lower-resource environments but is also capable of maximizing RD performance across a broad spectrum of resource configurations.
Fig. 8 presents the experimental results on the VVC Class B dataset. Most of the analyses observed on the UVG dataset, as discussed earlier, also apply here. However, due to the inherent characteristics of current INRs, which are more suitable for videos with fewer dynamic scenes, the overall gain relative to x265 (veryslow) on this dataset is lower than that observed on the UVG dataset. Nevertheless, at larger $p$ settings, the RD performance of HiNeRV (ours) still surpasses that of x265 (veryslow). In addition, we present the BD-Rate results for both PSNR and MS-SSIM metrics in Table I to show the performance of HiNeRV (ours) over four bitrate points. Among all INR-based methods, HiNeRV (ours) achieves the best performance. For the PSNR metric, it outperforms x265 (veryslow) in half of the experimental setups. In terms of MS-SSIM, HiNeRV (ours) exceeds x265 (veryslow) in most setups, achieving up to a 50% bitrate reduction with maintaining equivalent MS-SSIM when $p=120$. To further illustrate the effectiveness of our method, Fig. 9 presents visualized comparisons between HiNeRV (base) and HiNeRV (ours). The results clearly show that HiNeRV (ours) restores finer details while requiring a lower bitrate. Furthermore, HiNeRV (base) is more prone to amplifying certain details, causing them to become more prominent and interfering with the background. In contrast, HiNeRV (ours) models these details more effectively, preventing them from disrupting background consistency.
IV-C Advantage of Different GOP Settings
@@
Memory Usage and Computational Complexity are critical considerations for practical deployment. Memory usage and computational complexity are critical considerations for practical deployment. Table II presents the memory consumption of HNeRV under different batch sizes and model sizes. We observe that as the batch size (bs) and model size increase, memory usage also rises. For a 10 MB HiNeRV model, a batch size of 4 approaches the limit on a consumer-grade GPU with 24 GB of VRAM. While HNeRV is already a memory-efficient INR model due to its balanced parameter distribution, other INR models tend to consume even more memory. In our experiments, we also tested HiNeRV, which exhibits significantly higher memory usage, necessitating the use of GPUs with larger memory capacities for certain experiments.
Given the diverse range of encoding tasks across various scenarios, the compression model must operate on devices with different hardware configurations. In our framework, small GOP settings only require fitting smaller models, providing users with the flexibility to adjust hyperparameters according to their needs. More generally, our framework captures temporal correlations across multiple GOPs through autoregression, allowing each GOP to be loaded separately onto the GPU for training while maintaining an efficient compression ratio.
Fitting Speed influence not only deployment by also training efficiency. Apart from low GPU memory consumption, another advantage of a small model is the reduction in forward-pass computation time, which impacts both training and inference speeds. Table II presents the time required for a single training iteration of one frame (one forward and backward pass), showing that smaller model sizes result in faster processing speeds. Consequently, smaller $p$ settings ebables faster training. Early in training (not converged), the PSNR is higher for small $p$ with the same training time. In addition, the reduced computational cost during inference allows for operation on more lightweight devices, also making the approach suitable for Single-Encoding Multiple-Decoding (SEMD) scenario.
TABLE II: GPU memory cost (MB), Macs and training time cost per image(once forward+backward) with different $p$ (GOP) setting. The ”-” stands cuda out of memory.
ModelSize 0.3MB 1.0MB 3.0MB 5.0MB 10.0MB
Mem(bs1) 2500 2832 4014 4244 6062
Mem(bs2) 4326 4966 6852 8534 15770
Mem(bs4) 8264 10258 13194 15386 24187
Mem(bs8) 16248 20114 - - -
Macs 11.13G 43.16G 134.53G 217.32G 419.85G
Train time
Per iteration
37.78ms 53.33ms 86.67ms 116.67ms 242.22ms
IV-D Ablation Study
TABLE III: Ablation study of UAR-NVC. Values in the table mean bpp/PSNR (dB).
@@ IV-D Ablation Study
To evaluate the effectiveness of each proposed module, we conducted ablation experiments. We tested two components of UAR-NVC: the II module and the QEC module. For the II module, we compared it against random initialization (R-init), which does not utilize any information from the trained parameters of previous models, and direct initialization (D-init), which directly transfers the trained parameters of the previous model as the initialization for the current model. For the RQEC module, we assessed performance in the absence of reference-based compression (w/o ref).
For $p=6/30$, we compare V0 and V3, as shown in Table III, and found that the reference compression design in the QEC module reduces bitrate by 12.7%/18.9% while improving the PSNR metric by 0.3 dB/0.224 dB. This improvement occurs because the reference compression setting captures the correlation between adjacent GOPs, further enhancing reconstruction quality through the RD optimization mechanism of the QEC module. Comparing V1 and V3, our II-init module provides only a marginal improvement. This is because, after approaching convergence, II-init and R-init have minimal impact on the final RD performance without reference.

### v2

@@ IV-A2 Implementation Details
To achieve different bitrate points, we follow previous work to adjust the model size [16, 24, 25]. For small $p$, the model sizes are (0.3, 0.45, 0.6, 0.9) M. For medium $p$, the model sizes are (0.6, 0.9, 1.8, 2.7) M. For large $p$, the model sizes are (3.0, 4.5, 6.0, 9.0) M. In addition, we set the rate-distortion trade-off parameter $\lambda$ to 5.0 for NeRV and HiNeRV, and 0.5 for HNeRV, as they have very different network architectures. We obtained these values using a fitting method similar to $\varepsilon$ described in Sec. III-E. The small, medium and large $p$ settings correspond to different learning rates and training epochs for NeRV and HNeRV: 5e-3/5e-3/2e-3 and 3k/1.5k/1.5k for I-model, and 2k/2k/1k for P-model. For HiNeRV, the learning rates are adjusted to 2e-3/2e-3/1e-3. The papramter selection follows the setting in their original paper [25, 24]. Besides, we train for more epochs when $p$ is smaller, as fewer data points per epoch require longer training.
@@
[…]
@@
[…]
@@ IV-A3 Evaluation Metrics
Figure 9: Rate Distortion performance on Bunny dataset with different $p$ (GOP).
TABLE II: Rate Distortion performance on different sequence of UVG.
GOP GOP6 GOP30 GOP120
Methods HiNeRV (base) HiNeRV (ours) HiNeRV (base) HiNeRV (ours) HiNeRV (base) HiNeRV (ours)
Metrics bpp PSNR(dB) bpp PSNR(dB) bpp PSNR(dB) bpp PSNR(dB) bpp PSNR(dB) bpp PSNR(dB)
HoneyBee 0.435 38.697 0.104 39.136 0.246 39.016 0.043 39.318 0.2 39.509 0.036 39.517
Bosphorus 0.434 41.083 0.231 42.229 0.245 40.998 0.136 41.884 0.2 41 0.139 42.07
Beauty 0.431 35.265 0.362 35.94 0.242 34.784 0.186 35.206 0.198 34.657 0.177 35.199
YachtRide 0.433 37.468 0.348 39.443 0.244 36.104 0.22 38.527 0.199 35.587 0.208 38.856
Jockey 0.432 39.033 0.282 39.652 0.243 38.788 0.158 39.343 0.198 38.535 0.128 39.233
ReadySetGo 0.433 37.096 0.346 38.915 0.244 36.798 0.208 39.151 0.198 36.885 0.17 39.441
Avg. 0.433 38.107 0.279 39.219 0.244 37.748 0.158 38.905 0.199 37.695 0.143 39.053
IV-B Main Results on Rate-Distortion Performance
Overall performance with UVG dataset. To evaluate the performance of our framework under different settings, we first conduct experiments on the UVG dataset. As shown in Fig. 7, both NeRV, HNeRV and HiNeRV exhibit significant improvements when integrated with our proposed modules. Specifically, when using NeRV (base), HNeRV (base) and HiNeRV (base) as anchors, and comparing the RD curve positions of NeRV (ours), HNeRV (ours) and HiNeRV (ours), we observe that the primary gain in HNeRV stems from bitrate savings, while the primary gain in HiNeRV arises from improved image quality. At the same time, the structure of NeRV is more ordinary, improving a few in both dimensions. This improvement is attributed to HiNeRV’s more complex grid design, which enables it to capture finer details and fully leverage the potential of a more advanced entropy model with flexible bit allocation. Conversely, HNeRV’s architecture is optimized for efficiency, featuring a compact structure that reduces memory consumption and accelerates inference speed. This streamlined design also imposes a limit on its maximum reconstruction quality, even as it achieves substantial bitrate reductions with our two modules. When evaluated using the MS-SSIM metric, both models exhibit a trend similar to the analysis in PSNR. However, compared to x265 (veryslow) as the anchor, INR-based methods achieve significantly better in MS-SSIM relative to PSNR. This discrepancy arises because INR models encode video data in a more holistic manner, leading to superior structural fidelity. In contrast, traditional block-based encoders struggle to effectively capture the structural relationships between blocks.
To further illustrate the effectiveness of our method, Fig. 10 presents visualized comparisons between HiNeRV (base) and HiNeRV (ours). The results clearly show that HiNeRV (ours) restores finer details while requiring a lower bitrate. Furthermore, HiNeRV (base) is more prone to amplifying certain details, causing them to become more prominent and interfering with the background. In contrast, HiNeRV (ours) models these details more effectively, preventing them from disrupting background consistency.
Performance across difference $p$ settings. Examining the performance gap across different $p$ settings, we observe that in smaller $p$ settings, the RD performance of INR-based model is inferior to that in larger $p$ settings. This is primarily because models with larger $p$ settings are more effective at reducing temporal redundancy. Consequently, choosing an appropriate $p$ value involves a trade-off dictated by resource constraints. In addition, we note that when $p=6$, HiNeRV (ours) performs less effectively than x265 (veryslow) at low bitrates but surpasses it at higher bitrates. This occurs because, when fitted with a smaller model, the HiNeRV structure has limited capacity to capture fine details, reducing its fitting ability. However, higher bitrate models alleviate this limitation, leading to improved overall performance. As $p$ increases, the performance of our method improves steadily and eventually surpasses that of x265 (veryslow), demonstrating the robust modeling capabilities of implicit representation models. Notably, $p=30$ represents a highly practical configuration, with encoding delays that remain acceptable in many real-world scenarios. In this setup, our RD performance exceeds that of x265 (veryslow), further validating the effectiveness of our framework. When $p=120$, HiNeRV (ours) significantly outperforms x265 (veryslow). This result demonstrates that our framework is not only suitable for lower-resource environments but is also capable of maximizing RD performance across a broad spectrum of resource configurations.
Overall performance with VVC Class B dataset. Fig. 8 presents the experimental results on the VVC Class B dataset. Most of the analyses observed on the UVG dataset, as discussed earlier, also apply here. However, due to the inherent characteristics of current INRs, which are more suitable for videos with fewer dynamic scenes, the overall gain relative to x265 (veryslow) on this dataset is lower than that observed on the UVG dataset. Nevertheless, at larger $p$ settings, the RD performance of HiNeRV (ours) still surpasses that of x265 (veryslow).
In addition, we present the BD-Rate results for both PSNR and MS-SSIM metrics in Table I to show the performance of HiNeRV (ours) over four bitrate points. Our approach has significant RD performance improvements across all base models. Among all INR-based methods, HiNeRV (ours) achieves the best performance. For the PSNR metric, it outperforms x265 (veryslow) in half of the experimental setups. In terms of MS-SSIM, HiNeRV (ours) exceeds x265 (veryslow) in most setups, achieving up to a 50% bitrate reduction with maintaining equivalent MS-SSIM when $p=120$.
In our framework, the dependency between different GOPs within the same GOM is relatively weak. Decoding a frame in the current GOP does not require the explicit prior decoding of frames from other dependent GOPs. As the resulting random-access latency is comparable to that of x264/x265 with a GOP size of p, we chose that setting for our primary comparison. We also calculated the BD-Rate for PSNR against x265 (veryslow) with a GOP size of (p x m) on the UVG dataset. For HiNeRV (ours)-GOP6, the BD-rate relative to x265 changed from 20.43% (vs. GOP 6) to 42.91% (vs. GOP 6x5=30). For HiNeRV (ours)-GOP30, the BD-rate relative to x265 changed from -17.60% (vs. GOP 30) to -15.51% (vs. GOP 30x5=150). Although the performance margin is reduced under this comparison setting, the results for HiNeRV (ours) remain highly competitive, especially for longer GOPs.
@@ IV-B Main Results on Rate-Distortion Performance
IV-C Performance on More Dimensions
Performance with difference video content. To validate the effectiveness of our framework on diverse video content, we present the RD performance of HiNeRV (base) and HiNeRV (ours) on various sequences from the UVG dataset. These include sequences with limited motion, such as “HoneyBee”, and those with significant motion, like “ReadySetGo”. According to the results in Table II, our framework demonstrates a significant performance advantage across all sequences. Notably, our RQEC module can dynamically adjust the optimization direction based on the RD behavior during training. For “HoneyBee”, a sequence where improving reconstruction quality is challenging, HiNeRV (ours) achieves substantial bitrate reduction. Conversely, for sequences with more room for quality improvement, such as “ReadySetGo”, HiNeRV (ours) significantly enhances reconstruction quality.
Performance with difference video resolution (Bunny dataset). We also evaluated our framework’s effectiveness on the Bunny dataset, with results presented in Fig. 9. It is observed that HNeRV demonstrates better adaptability to such low-resolution video sequences, especially when $p$ is small. However, despite the inherent model architectures sometimes leading to performance characteristics that may not perfectly align with typical expectations across different configurations, our module consistently enhances the RD performance.
Compare with other methods. Having validated the flexibility of our framework and the effectiveness of its modules, we compare its performance against DCVC [13] and NIRVANA [27]. For DCVC, we calculated the BD-Rate of HiNeRV (ours) relative to DCVC under various $p$ settings, as presented in Table III. Although our RD performance is comparatively lower at smaller $p$ values, it progressively surpasses DCVC as $p$ increases. Owing to our framework’s flexibility, users can select different configurations based on their practical requirements and resource constraints. Regarding NIRVANA, we set $p$ to 3 and compare with the results it provided in the appendix on UVG (1920x1080), as no source code is provided. As shown in Table IV. While NIRVANA’s original paper reported slightly superior RD performance over the original NeRV, our NeRV (ours) manages to outperform NIRVANA, even at this $p=3$ setting. Furthermore, HiNeRV (ours) demonstrates substantially superior RD performance. Despite NIRVANA employing a lightweight network architecture, our framework supports various NeRV models, allowing users to choose models optimized for speed or RD performance, without being restricted by a specific model structure.
TABLE III: The BD-rate (%) of HiNeRV (ours) over DCVC, measured in PSNR.
GOP6/5 GOP30/25 GOP120/100
UVG 22.233 -15.731 -46.154
ClassB 80.427 25.169 -20.093
TABLE IV: Detailed Comparison with NIRVANA over UVG.
NIRVANA
NeRV (ours)
GOP3-1.2M
HiNeRV (ours)
GOP3-1.2M
PSNR bpp RSNR bpp RSNR bpp
HoneyBee 38.83 0.51 39.624 0.52 39.907 0.366
Bosphorus 40.53 0.68 41.24 0.614 43.353 0.496
Beauty 35.77 0.96 35.576 0.646 37.577 0.974
YachtRide 37.94 1.03 37.339 0.761 41.741 0.762
Jockey 37.56 0.96 39.103 0.636 41.027 0.71
ReadySetGo 35.43 1.26 35.234 0.782 41.255 0.745
Avg. 37.677 0.9 38.019 0.66 40.81 0.676
IV-D Advantage of Different GOP Settings
@@
Memory Usage and Computational Complexity. Table V presents the memory consumption of HiNeRV under different model sizes. We observe that as the model size increase, memory usage also rises. While the model sizes is lager than 9.0M, the training of HiNeRV could not be employed in an RTX 3090 GPU anymore. More generally, our framework captures temporal correlations across multiple GOPs through autoregression, allowing each GOP to be loaded separately onto the GPU for training while maintaining an efficient compression ratio.
Fitting Speed. Apart from low GPU memory consumption, another advantage of a small model is the reduction in forward-pass computation time, which impacts both training and inference speeds. Table V presents the time required for training one frame (one forward and backward pass), showing that smaller model sizes result in faster processing speeds. Consequently, smaller $p$ settings ebables faster training. Early in training (not converged), the PSNR is higher for small $p$ with the same training time. In addition, the reduced computational cost during inference allows for operation on more lightweight devices, also making the approach suitable for Single-Encoding Multiple-Decoding (SEMD) scenario.
TABLE V: Detailed Memory Consumption and Speed Comparison across Base Models, Modules, and Model Sizes.
Model
Name
Model
Size
Module Macs Encode
Entropy Coding of
Model Parameters
Decode
Memory
Time per
Frame (s)
FPS
Encode
Time (s)
Decode
Time (s)
Memory
Time per
Frame (s)
FPS
NeRV 1.8M RQEC 124.32G 3508.67 0.07008 14.27 0.1058 0.23411 1478.0 0.01336 74.85
NeRV 1.8M QAT 124.32G 3464.0 0.0679 14.73 0.46824 0.45308 1505.33 0.01316 75.99
HNeRV 1.8M RQEC 67.48G 2966.0 0.06091 16.42 0.11072 0.23413 1596.0 0.00766 130.55
HNeRV 1.8M QAT 67.48G 2942.0 0.05932 16.86 0.55829 0.44402 1412.0 0.00743 134.59
HiNeRV 1.8M RQEC 125.55G 11196.67 0.22136 4.52 0.11698 0.27063 2923.33 0.08826 11.33
HiNeRV 1.8M QAT 125.55G 11020.0 0.21089 4.74 0.94722 0.48933 2826.67 0.08755 11.42
HiNeRV 0.3M RQEC 21.19G 5470.0 0.17695 5.65 0.02995 0.08386 1415.33 0.07437 13.45
HiNeRV 0.3M QAT 21.19G 5370.67 0.16828 5.94 0.6322 0.10451 1397.33 0.07371 13.57
HiNeRV 3.0M RQEC 208.24G 13838.0 0.29304 3.41 0.19746 0.43358 5102.67 0.11662 8.57
HiNeRV 3.0M QAT 208.24G 13698.67 0.28437 3.52 1.13121 0.85308 4480.0 0.11652 8.58
HiNeRV 6.0M RQEC 413.21G 19298.67 0.35207 2.84 0.40545 0.84315 9161.33 0.13703 7.3
HiNeRV 6.0M QAT 413.21G 19244.67 0.34168 2.93 1.84024 1.56098 8287.33 0.13649 7.33
HiNeRV 9.0M RQEC 613.12G 23596.67 0.50683 1.97 0.66516 1.4059 12263.33 0.18056 5.54
HiNeRV 9.0M QAT 613.12G 23338.0 0.49542 2.02 2.79206 2.37955 11045.33 0.18061 5.54
IV-E Ablation Study
TABLE VI: Ablation study of UAR-NVC. Values in the table mean bpp/PSNR (dB).
@@ IV-E Ablation Study
To evaluate the effectiveness of each proposed module, we conducted ablation experiments. We tested two components of UAR-NVC: the II module and the RQEC module. For the II module, we compared it against random initialization (R-init), which does not utilize any information from the trained parameters of previous models, and direct initialization (D-init), which directly transfers the trained parameters of the previous model as the initialization for the current model. For the RQEC module, we assessed performance in the absence of reference-based compression (w/o ref).
For $p=6/30$, we compare V0 and V3, as shown in Table VI, and found that the reference compression design in the RQEC module reduces bitrate by 12.7%/18.9% while improving the PSNR metric by 0.3 dB/0.224 dB. This improvement occurs because the reference compression setting captures the correlation between adjacent GOPs, further enhancing reconstruction quality through the RD optimization mechanism of the RQEC module. Comparing V1 and V3, our II-init module provides only a marginal improvement. This is because, after approaching convergence, II-init and R-init have minimal impact on the final RD performance without reference.
@@
To evaluate potential overhead introduced by our RQEC module, we split the whole process into encode (training), entropy coding and decode (inference), reporting training/inference speed, GPU memory usage, computational complexity, and entropy coding time under various settings in Table V. We observe that, The RQEC module adds approximately 10% to GPU memory usage during inference, which is an acceptable trade-off. For entropy coding, RQEC is markedly faster, which may be attributed to the use of different entropy coding libraries (the entropy coding library used by QAT in HiNeRV [^1] https://pypi.org/project/torchac/ does not support the RQEC module well, so we used the constriction library [^2] https://pypi.org/project/constriction/ instead). These findings are consistent across various base models and sizes, demonstrating that the RQEC module does not introduce substantial overhead.